OpenStudy (anonymous):

If you perform 10 joules of work lifting a 10-N box from the floor to a shelf, how high is the shelf?

OpenStudy (anonymous):

mgh= energy / work mgh=10 (10) h=10 h=1

OpenStudy (fellowroot):

I think that the box needs to be in kg not in Newtons

OpenStudy (ghazi):

mg is the weight of box and that is equal to 10N, so as she has done above m*g*h= P.E 10*h= 10 h=1 m height is 1m